# Test-Time Training Scaling for Chemical Exploration in Drug Design

### v1

@@ Abstract
Chemical language models for molecular design have the potential to find solutions to multi-parameter optimization problems in drug discovery via reinforcement learning (RL). A key requirement to achieve this is the capacity to ”search” chemical space to identify all molecules of interest. Here, we propose a challenging new benchmark to discover dissimilar molecules that possess similar bioactivity, a common scenario in drug discovery, but a hard problem to optimize. We show that a population of RL agents can solve the benchmark, while a single agent cannot. We also find that cooperative strategies are not significantly better than independent agents. Moreover, the performance on the benchmark scales log-linearly with the number of independent agents, showing a test-time training scaling law for chemical language models.
Machine Learning, ICML
[…]
@@ 1 Introduction
Identifying better, safer, and more efficacious drug candidates faster can save hundreds of millions of dollars (Bender & Cortés-Ciriano, 2021), and bring medicines to patients years quicker. However, the vastness of chemical space (Polishchuk et al., 2013) is a central challenge. It is intractable to computationally enumerate all drug-like molecules, let alone evaluate them, synthesize them and experimentally test them in the lab. Traditionally, virtual screening is conducted of virtual libraries that contain just tiny fractions of drug-like chemistry that can be purchased readily and experimentally tested. On the other hand, goal-directed generative AI can generate novel chemical structures with optimal predicted properties while implicitly searching much larger chemical spaces (Arús-Pous et al., 2019a). Thus, harnessing the full potential of generative AI to facilitate drug design will have considerable positive impacts on drug discovery as a whole.
To fully exploit generative AI potential, achieving a thorough exploration of chemical space is critical. The more diverse solutions identified the more advantageous it is. Starting a drug discovery campaign from multiple diverse hits can offer backup series with alternative chemical and biological profiles, decreasing risk by increasing shots at goal (Brown & Boström, 2018). Additionally, probing novel areas of chemical space can avoid already patented compounds (Shimizu et al., 2023). Furthermore, in practice, many generative models undergo ”mode collapse” (a narrowing of chemical space) when optimizing for particular predicted properties, particularly those that are predicted by machine learning models with limited applicability domains arising from epistemic uncertainty (Preuer et al., 2018; Renz et al., 2019).
Reinforcement learning (RL) is a family of optimization algorithms that can train an agent (a decision-making entity) to maximize a reward within a Markov decision process (Sutton & Barto, 2018). In generative molecular design, the generative model acts as the RL agent that learns to adapt molecule generation to maximize a given reward. One or more predicted molecular properties can be combined into an ’oracle’ that provides the reward. A range of generative models as RL agents for molecular design have been demonstrated including heuristic rules combined with deep Q-networks (Zhou et al., 2019), deep language model policy functions (Olivecrona et al., 2017; Ståhl et al., 2019), or deep graph network policy functions (Atance et al., 2022). However, it is the deep language model policy functions, also referred to as chemical language models (CLMs) (Segler et al., 2018; Grisoni, 2023; Skinnider et al., 2021), that have emerged as the most popular generative drug design algorithms (Olivecrona et al., 2017; Thomas et al., 2022b; Guo & Schwaller, 2024), as they have demonstrated top performance on multiple benchmarks (Brown et al., 2019; Polykovskiy et al., 2020; Thomas et al., 2022a).
CLMs are autoregressive language models such as recurrent neural networks (RNNs) or GPT-style models trained on next token prediction typically using 1D SMILES (Weininger, 1988) strings that represent the 2D graph of a molecule. Seminal works (Olivecrona et al., 2017; Popova et al., 2018) applied RL to CLMs resulting in an agent capable of generating de novo molecules with desirable properties. The body of work now surrounding these models includes algorithmic variations to improve the learning efficiency (Bjerrum et al., 2023; Thomas et al., 2022b; Guo & Schwaller, 2024; Thomas et al., 2025), the practical utility (Fialková et al., 2021; Thomas et al., 2024a), as well as experimental validation of workflows utilizing such models for de novo drug design (Korshunova et al., 2022; Quinn et al., 2024; Thomas et al., 2024b). Interestingly, REINFORCE (Sutton & Barto, 2018) based RL algorithms remain superior for CLMs (Bou et al., 2024), likely due to strong policy initialization from pre-training (Ahmadian et al., 2024).
In the field of Large Language Models (LLMS), Test-Time Training (TTT) has emerged as a successful approach to improve abstract reasoning (Akyürek et al., 2024). TTT is a strategy that updates the parameters of a pre-trained LLM on a specific test-time task, where they are reset to the pre-trained parameters for the next task. Hence, the focus is on the predictions of the temporarily updated LLM instead of the updated LLM with its new parameters. Moreover, scaling inference-time has shown to improve performance on benchmarks log-linearly with the number of samples generated at inference (Brown et al., 2024). Together, they have led to landmark improvements in abstract reasoning benchmarks such as the ARC Challenge (Chollet, 2019).
Multi-agent reinforcement learning (MARL) involves the cooperation, competition, or evolution of multiple agents acting in an environment to enhance learning (Buşoniu et al., 2010). This is a proven effective strategy to reach human or greater level performance in StarCraft (Vinyals et al., 2019) and capture the flag (Jaderberg et al., 2019), or to learn cooperative microeconomic strategies (Johanson et al., 2022). Most relevant to the work presented here is the capacity of a population of agents to increase the diversity of solutions via cooperative learning by enforcing divergent behavior between agents (Parker-Holder et al., 2020). Although MARL is effective in fields such as game theory, few examples exist applying MARL to generative molecular design.
In this work, we provide two central contributions:
  1. 1.
    We introduce a new hard de novo drug design benchmark focused on efficient Molecule Exploration (MolExp). Unlike existing benchmarks, MolExp emphasizes discovering structurally diverse molecules that retain similar bioactivity, acting as an exploration challenge where multiple distinct regions of chemical space get high rewards and all should be found.
  2. 2.
    We show that a population of independent agents is better at chemical space exploration on the benchmark than a single agent trained for a long time, even with exploration bonuses. We also show that performance on the benchmark scales with the number of agents in the population. Lastly, we investigate several cooperative strategies between multiple agents to try to enhance exploration efficiency.
@@ 2 Related Work
Balancing the exploration-exploitation trade-off is a constant challenge in RL, including RL for drug design (Langevin et al., 2024). Exploitation initially maximizes the reward quicker but may lead to a lower overall reward, while exploration may lead to higher overall reward and increased diversity of solutions but takes longer. Diversity has consistently been an important topic in generative molecular design and is increasingly included in the evaluation of performance (Thomas et al., 2022a; Renz et al., 2024). Several approaches have been proposed to increase diversity, for example, using dual networks to balance exploitation-exploration during sampling (Liu et al., 2019), penalizing the reward of molecules similar to previously generated ones (Blaschke et al., 2020), or increasing the reward of molecules that are dissimilar to those previously generated (Thiede et al., 2022). Recent work shows that random network distillation (RND) (Burda et al., 2018) is particularly effective at increasing molecular diversity with low overhead due to not requiring an explicit memory of previously generated molecules (Svensson et al., 2024; Thomas et al., 2025).
Previous work has applied multi-agent RL to de novo drug design. Hu et al. (2024) deployed up to 4 GPT agents as CLMs that use the REINVENT loss formulation with an additional penalty to discourage the $k$th agent from being similar to previous agents. This multi-agent setup outperformed baseline algorithms on the GuacaMol benchmark (Brown et al., 2019). However, this benchmark consists of mostly simple tasks (as also stated by the authors). As demonstrated by their experiment comparing 1, 2, 4 and 8 agents all having near maximum performance and all within error. Moreover, independent agents and other cooperative strategies were not investigated. Therefore it remains to be seen whether MARL improves performance significantly, by how much, and which cooperative strategies work best.
Generative molecular design benchmarks must reflect realistic scenarios in drug design for practical application, as gold standard evaluation via synthesis and experimental validation of molecules is too costly at scale. The GuacaMol benchmark (Brown et al., 2019) proposed a suite of 20 tasks, the authors highlighted that $\sim 15/20$ tasks are easily solved by the generative algorithms. The MolOpt benchmark limited these tasks to a budget of 10,000 molecules during optimization, shifting the focus to efficiency, however, this likely over-rewards exploitative algorithms as opposed to explorative ones. New criteria have been defined to additionally measure the quality of chemistry generated which has changed rank performance (Thomas et al., 2022a; Renz et al., 2024), moreover, rank performance on this MolOpt benchmark didn’t reflect performance on more realistic optimization objectives such as predicted binding affinity via docking (Bou et al., 2024). Hence, there remains significant scope for new exploration-focused objectives that better reflect realistic optimization objectives in drug design.
@@ 3.1 Problem definition
The generative molecular design problem is the generation of molecules that are optimal in one or more properties defined by the drug discovery context, for example, drug-likeness and predicting binding affinity against a protein implicated in a disease. The scoring function (also referred to as an oracle) $s(x)\in[0,1]$ may combine one or more property predictors $p_{i}(x)$ that score a given molecule $x$ returning a real value $p_{i}:X\rightarrow\mathbb{R}$. Multi-parameter optimization can be achieved by combining multiple predicted properties, the simplest example being to take the arithmetic mean $\sum_{i}t_{i}(p_{i}(x))$ ensuring that $t(p_{i}(x))\in[0,1]$ by using a transformation function $t_{i}$. Overall, the final score reflects the desirability of the molecule for the given context, however, optimizing a score combining many properties is usually non-trivial as improving one property can be detrimental to another resulting in a Pareto-optimization problem.
$\max_{x\in X}s(x)$ (1)
[…]
@@
We propose that a generative model benchmark should adequately mimic this concept of identifying multiple, seemingly unrelated molecular targets in chemical space (Figure 1, right). Firstly, previously published benchmarks typically don’t capture this with objectives focused on a single narrow target, or a single broad target (Figure 1, left). These do not adequately test the ability of an algorithm to explore chemical space if it is already receiving high rewards. Secondly, there are post-hoc metrics that measure the diversity of solutions (Thomas et al., 2022a; Renz et al., 2024), however, this only measures the diversity inherently embedded in the objective. Therefore, designing an objective to contain multiple distinct targets provides a greater opportunity to observe inherent exploration capability. Lastly, previous benchmarks predominantly measure performance on the maximum $S(x)$ achieved, whereas, we propose to measure performance on maximum $p_{i}(x)$ for each molecular target. This focuses evaluation on the algorithm’s ability to explore all high-rewarding regions.
@@ 3.2 Reinforcement learning with chemical language models
CLMs utilize a language representation, commonly SMILES (Weininger, 1988), to leverage natural language models such as recurrent neural networks (RNNs) or GPT-style models. For example, SMILES represent the molecular structure of paracetamol with the string ”CC(=O)Nc1ccc(O)cc1”. These models operate autoregressively, predicting the next token $x_{t}$ in a sequence $X=(x_{1},x_{2},\ldots,x_{T})$ based on the context provided by previous tokens. The process can be formalized as maximizing the conditional probability of each token given the preceding sequence $P(X)=\prod_{t=1}^{T}P(x_{t}|x_{<t})$, where $x_{<t}$ denotes the tokens up to $t-1$. The model is trained using the cross-entropy loss function which is equivalent to the negative log-likelihood Equation 3. Where $T_{i}$ is the length of the $i$-th sequence, and $\theta$ are the model parameters.
@@
We deviate to vanilla REINFORCE by reshaping the reward to regularize the agent policy to the prior policy $\pi_{prior}$ with a coefficient $\sigma$ and by adding an exponent $\alpha$, as proposed by Thomas et al. (2025). These components improve the quality of molecules proposed by the agent and improve performance by steepening the gradients in the reshaped reward landscape. In addition to reshaping the reward, we implement experience replay to augment on-policy data with off-policy data at each iteration. A replay buffer is kept of high-reward molecules which is sampled with prioritization proportional to their reward. This additionally improves learning efficiency (Gao et al., 2022; Thomas et al., 2025).
@@
3.3 Multi-agent reinforcement learning
MARL extends the framework of single-agent RL to systems with multiple independent or interacting agents, each learning to optimize its own policy in an environment. The problem is the same as defined before, but now with $\mathcal{N}$ agents.
@@ 3.3 Multi-agent reinforcement learning
In the interaction case, we focused on methods of cooperation within a population-learning setting which typically results in more diverse solutions (Oroojlooy & Hajinezhad, 2023). In this case, using $\langle\mathcal{S},\mathcal{A}^{i},R,P^{i}\rangle$ for $i\in\mathcal{N}$ where the states $S$ and $R$ may be shared across agents in a cooperative manner. For simplicity, here we refer to the complete molecule and final state $s_{T}$ simply as the state. We investigated the following methods of cooperation (for additional detail and equations see Appendix B):
@@
Re-implementation of the cooperative strategy applied by Hu et al. (2024) which seeks to maximize the log-probability difference of on-policy states between the current agent and other agents.
@@
Diversity via Determinant (DvD) inspired by Parker Holder et al. (2020). Calculate a policy behavior embedding and maximize the divergence in policy behavior as measured by the determinant of the behavioral embedding transformed by a kernel function.
@@ 4 Experiments
To assess the ability of single and multiple RL agents to explore chemical space, we conducted three experiments.
@@
    We defined a molecule exploration benchmark and evaluated baseline algorithm performance.
@@
    We tested the effectiveness of different population sizes of multiple independent agents, and compared them to equivalent single agents.
@@
    We investigated different strategies of cooperation within a population of agents to enhance exploration efficiency.
@@ 4.1 Molecule exploration benchmark
To mimic our proposed problem formulation of identifying all hits in chemical space requiring intrinsic exploration, we designed a new molecular exploration (MolExp) benchmark. To embed this benchmark in real-world discovery, it was designed based on the converse of the similarity principle  [^1] In drug discovery, the similarity principle states that similar molecules $Sim(m_{1},m_{2})\geq\epsilon$ possess similar bioactivity $|f(m_{1})-f(m_{2})|\leq\delta$. However, this does not prove the converse and many exceptions exist that are more difficult to model., i.e., to identify multiple dissimilar molecules that exhibit similar desirable properties such as bioactivity. Note that activity cliffs are another example converse to the similarity principle that have been used to demonstrate more challenging scenarios for predictive models (Kwapien et al., 2022; Van Tilborg et al., 2022).
@@
Each run is replicated on the benchmark 5 times and as with previous benchmarks (Gao et al., 2022) a default budget of 10,000 molecules is used.
@@
Oracles The benchmark consists of four tasks, each using an oracle to evaluate molecules. These are based on drug candidates and their biological targets, including anti-psychotic drugs (AP), Adenosine A2A (A2A) receptor ligands, beta-secretase 1 (BACE1) inhibitors, and epidermal growth factor inhibitors (EGFR). These tasks contain 2, 3, 3, & 4 objectives respectively, where the objective is the similarity to a target molecule $sim(m_{i},m_{t})$. For each task, only a single reward is provided based on the similarity to the nearest target molecule $R(m)=max(sim(m_{i},m_{t_{1}}),...,sim(m_{i},m_{t_{N}}))$. Due to CLMs acting in string space, we chose a string-edit similarity function. This, reflects chemical space search scenarios where distinct regions of reward signal exist. A more detailed description of the molecular targets chosen and the similarity function can be found in Appendix C.
@@
Metric Performance is measured by calculating the product of the top N most similar molecules to each target molecule respectively $\prod_{t\in T}\frac{\sum_{i\in N}sim(m_{i},m_{t})}{N}$. This metric focuses on the ability generate similar molecules to all objectives. Unless stated otherwise, we use $N=1$.
@@
Baselines We tested virtual screening and multiple RL algorithms (see Appendix D) on their inherent ability for chemical exploration on the MolExp benchmark. Table 1 shows that the recently published ACEGENMolOpt (Thomas et al., 2025), performed best. Figure 13 illustrates that none of the tested algorithms optimize all molecular targets within the default budget (for a single replicate). Moreover, that the best scores are due to improved efficiency in maximizing similarity to a single target molecule. Therefore, there is still considerable room for improvement.
@@
Estimated max An estimated maximum performance can be calculated for the chosen agent RL configuration (ACEGENMolOpt) by training the agent to maximize the similarity to each individual molecular target with the reward $R(m)=sim(m_{i},m_{t})$ only, Figure 14. This also tests the agents ability to optimize for the similarity to each individual molecular target in the first place which is an expected prerequisite for performance on the benchmark. Averaging the Top-1 similarity achieved over the 5 replicates and calculating the product over the group of molecular target objectives in each task, and then summing over all tasks led to an estimated achievable maximum benchmark score of 2.91.
Table 1: Baseline performance of virtual screening and policy-gradient based RL algorithms on the MolExp benchmark within a budget of 10,000 molecules. ACEGENMolOpt outperforms all other methods, while virtual screening outperforms two baselines algorithms vanilla REINFORCE and REINVENT.
@@
Task SCREEN REINFORCE REINVENT REINVENTMolOpt AHC ACEGENPractical ACEGENMolOpt
AP 0.53 ± 0.04 0.56 ± 0.05 0.57 ± 0.04 0.63 ± 0.02 0.62 ± 0.06 0.64 ± 0.02 0.64 ± 0.04
A2A 0.38 ± 0.03 0.36 ± 0.02 0.39 ± 0.08 0.38 ± 0.03 0.40 ± 0.02 0.40 ± 0.07 0.44 ± 0.06
BACE1 0.29 ± 0.08 0.23 ± 0.02 0.21 ± 0.01 0.27 ± 0.03 0.23 ± 0.01 0.23 ± 0.03 0.28 ± 0.03
EGFR 0.17 ± 0.04 0.15 ± 0.02 0.16 ± 0.02 0.22 ± 0.03 0.17 ± 0.03 0.17 ± 0.02 0.25 ± 0.05
Sum 1.36 ± 0.10 1.30 ± 0.06 1.33 ± 0.09 1.50 ± 0.06 1.41 ± 0.07 1.44 ± 0.08 1.62 ± 0.09
4.2 Test-time training with RL for chemical exploration

### v2

# Test-Time Training Scaling Laws for Chemical Exploration in Drug Design

@@ Abstract
Chemical Language Models (CLMs) leveraging reinforcement learning (RL) have shown promise in de novo molecular design, yet often suffer from mode collapse, limiting their exploration capabilities. Inspired by Test-Time Training (TTT) in large language models, we propose scaling TTT for CLMs to enhance chemical space exploration. We introduce MolExp, a novel benchmark emphasizing the discovery of structurally diverse molecules with similar bioactivity, simulating real-world drug design challenges. Our results demonstrate that scaling TTT by increasing the number of independent RL agents follows a log-linear scaling law, significantly improving exploration efficiency as measured by MolExp. In contrast, increasing TTT training time yields diminishing returns, even with exploration bonuses. We further evaluate cooperative RL strategies to enhance exploration efficiency. These findings provide a scalable framework for generative molecular design, offering insights into optimizing AI-driven drug discovery.
@@ 1 Introduction
The vast chemical space, estimated to contain over $10^{60}$ drug-like molecules [1], presents a formidable challenge in drug discovery. Efficiently identifying diverse, safe, and efficacious drug candidates can accelerate development timelines and reduce costs by hundreds of millions of dollars [2]. Traditional virtual screening of limited chemical libraries captures only a fraction of possible compounds, restricting exploration. In contrast, goal-directed generative AI, particularly Chemical Language Models (CLMs), enables de novo design of novel molecules with optimized properties, implicitly navigating expansive chemical spaces [3].
Thorough exploration of chemical space is critical to maximize the potential of generative AI in drug design. Identifying structurally diverse molecules with similar bioactivity increases the likelihood of discovering backup series with distinct chemical and biological profiles, mitigating risks in drug development [4]. Moreover, exploring novel chemical regions can circumvent patented compounds [5]. However, many generative models, including CLMs, suffer from mode collapse when optimizing for predicted properties, particularly those reliant on machine learning models with limited applicability domains due to epistemic uncertainty [6, 7].
CLMs, autoregressive models trained on SMILES strings [8], have become the leading approach for generative drug design [9, 10, 11], excelling in benchmarks like GuacaMol [12, 13]. By integrating reinforcement learning (RL), seminal works [14, 15] have enabled CLMs to generate molecules with tailored properties. Advances in RL algorithms [16, 17, 18, 19] and practical applications [20, 21] have further enhanced their utility, with REINFORCE-based methods remaining superior due to robust pre-training [22, 23, 24].
Inspired by Test-Time Training (TTT) in large language models, where model parameters are temporarily updated for specific tasks [25], we frame RL-based optimization of pre-trained CLMs as a form of TTT. In LLMs, scaling inference-time sampling has shown log-linear performance improvements [26]. Similarly, scaling TTT for CLMs—by increasing the number of RL agents or training time—offers potential to enhance chemical space exploration. Cooperative RL strategies further amplify exploration through population-based learning [27], though their application in molecular design remains underexplored.
In this work, we make two key contributions. First, we demonstrate that scaling TTT by increasing the number of independent RL agents follows a log-linear scaling law, significantly enhancing exploration efficiency. Conversely, extending TTT training time yields limited benefits, even with exploration bonuses or diversity filters. We also propose and evaluate cooperative RL strategies to further improve exploration efficiency. Second, we introduce MolExp, a novel benchmark to measure Molecule Exploration. MolExp requires the rediscovery of structurally diverse molecules with comparable bioactivity. Unlike existing benchmarks, MolExp requires exploration of all high-reward regions of chemical space, extending the boundaries of real-world drug design.
This work provides a robust framework for enhancing chemical space exploration, addressing critical challenges in AI-driven drug discovery. By leveraging TTT and cooperative RL, we offer practical tools through MolExp and scalable strategies that can accelerate the discovery of diverse drug candidates.
@@ 2 Related Work
Balancing the exploration-exploitation trade-off is a constant challenge in RL, including RL for drug design [28]. Exploitation initially maximizes the reward quicker but may lead to a lower overall reward, while exploration may lead to a higher overall reward and increased diversity of solutions but takes longer. Diversity has consistently been an important topic in generative molecular design and is increasingly included in the evaluation of performance [29, 30]. Several approaches have been proposed to increase diversity, for example, using dual networks to balance exploitation-exploration during sampling [31], penalizing the reward of molecules similar to previously generated ones [32], or increasing the reward of molecules that are dissimilar to those previously generated [33]. Recent work shows that random network distillation (RND) [34] is particularly effective at increasing molecular diversity with low overhead due to not requiring an explicit memory of previously generated molecules [35, 19].
Previous work has applied population-based RL to de novo drug design. Hu et al. [36] deployed up to 4 GPT agents as CLMs that use the REINVENT loss formulation with an additional penalty to discourage the $k$-th agent from being similar to previous agents. This multi-agent setup outperformed baseline algorithms on the GuacaMol benchmark [12]. However, this benchmark consists of mostly simple tasks (as also stated by the authors). As demonstrated by their experiment comparing 1, 2, 4 and 8 agents all having near maximum performance and all within standard deviation. Moreover, independent agents and other cooperative strategies were not investigated. Therefore it remains to be seen whether cooperative strategies improve performance significantly, by how much, and which cooperative strategies work best.
Generative molecular design benchmarks must reflect realistic scenarios in drug design for practical application, as gold standard evaluation via synthesis and experimental validation of molecules is too costly at scale. The GuacaMol benchmark [12] proposed a suite of 20 tasks, the authors highlighted that $\sim 15/20$ tasks are easily solved by the generative algorithms. The MolOpt benchmark limited these tasks to a budget of 10,000 molecules during optimization, shifting the focus to efficiency, however, this likely over-rewards exploitative algorithms as opposed to explorative ones. New criteria have been defined to additionally measure the quality of chemistry generated which has changed rank performance [29, 30], moreover, rank performance on this MolOpt benchmark didn’t reflect performance on more realistic optimization objectives such as predicted binding affinity via docking [23]. Hence, there remains significant scope for new exploration-focused objectives that better reflect realistic optimization objectives in drug design.
@@ 3.1 Problem definition
The generative molecular design problem is the generation of molecules that are optimal in one or more properties defined by the drug discovery context, for example, drug-likeness and predicting binding affinity against a protein implicated in a disease. The scoring function (also referred to as an oracle) $s(x)\in[0,1]$ may combine one or more property predictors $p_{i}(x)$ that score a given molecule $x$ returning a real value $p_{i}:X\rightarrow\mathbb{R}$. The final score reflects the desirability of the molecule for the given context, where the goal is to generate a molecule with the best score.
$\arg\max_{x\in X}s(x)$ (1)
However, chemical space is vast, property predictors are inherently inaccurate with high uncertainty, and Pareto optimal solutions have different property profiles. Hence, multiple molecules exist across chemical space that are considered hits if the score is above a given threshold $T$ (or sometimes more simply the top-$k$ molecules). We propose that the problem of generative molecular design should extend to finding all possible hits across chemical space, and postulate that this better aligns with practical requirements in drug discovery.
@@
We propose that a generative model benchmark should adequately mimic this concept of identifying multiple, even structurally unrelated molecular targets in chemical space (Figure 1, right). Firstly, previously published benchmarks typically don’t capture this with objectives focused on a single narrow target, or a single broad target (Figure 1, left). These do not adequately test the ability of an algorithm to explore chemical space if it is already receiving high rewards. Secondly, there are post-hoc metrics that measure the diversity of solutions [29, 30], however, this only measures the diversity inherently embedded in the objective. Therefore, designing an objective to contain multiple distinct targets provides a greater opportunity to observe inherent exploration capability. Lastly, previous benchmarks predominantly measure performance on the maximum $S(x)$ achieved, whereas, we propose to measure performance on maximum $p_{i}(x)$ for each molecular target representing a maxima in the reward landscape. This focuses evaluation on the algorithm’s ability to explore all high-rewarding regions.
@@ 3.2 Reinforcement learning with chemical language models
CLMs utilize a language representation, commonly SMILES [8], to leverage natural language models such as recurrent neural networks (RNNs) or GPT-style models. For example, SMILES represent the molecular structure of paracetamol with the string “CC(=O)Nc1ccc(O)cc1". These models operate autoregressively, predicting the next token $x_{t}$ in a sequence $X=(x_{1},x_{2},\ldots,x_{T})$ based on the context provided by previous tokens. The process can be formalized as maximizing the conditional probability of each token given the preceding sequence $P(X)=\prod_{t=1}^{T}P(x_{t}|x_{<t})$, where $x_{<t}$ denotes the tokens up to $t-1$. The model is trained using the cross-entropy loss function which is equivalent to the negative log-likelihood Equation 3. Where $T_{i}$ is the length of the $i$-th sequence, and $\theta$ are the model parameters.
@@
We deviate to vanilla REINFORCE by reshaping the reward to regularize the agent policy to the prior policy $\pi_{prior}$ with a coefficient $\sigma$ and by adding an exponent $\alpha$, as proposed by Thomas et al. [19]. These components improve the quality of molecules proposed by the agent and improve performance by steepening the gradients in the reshaped reward landscape. In addition to reshaping the reward, we implement experience replay to augment on-policy data with off-policy data at each iteration. A replay buffer is kept of high-reward molecules which is sampled with prioritization proportional to their reward. This additionally improves learning efficiency [37, 19].
@@
3.3 Cooperative reinforcement learning
Cooperative RL extends the framework of single-agent RL to systems with multiple independent or interacting agents, each learning to optimize its own policy in an environment. The problem is the same as defined before, but now with $\mathcal{N}$ agents.
@@ 3.3 Cooperative reinforcement learning
In the interaction case, we focused on methods of cooperation within a population-learning setting which typically results in more diverse solutions [38]. In this case, using $\langle\mathcal{S},\mathcal{A}^{i},R,P^{i}\rangle$ for $i\in\mathcal{N}$ where the states $S$ and $R$ may be shared across agents in a cooperative manner. For simplicity, here we refer to the complete molecule and final state $s_{T}$ simply as the state. We investigated the following methods of cooperation (for additional detail and equations see Appendix B):
@@
Re-implementation of the cooperative strategy applied by Hu et al. [36] which seeks to maximize the log-probability difference of on-policy states between the current agent and other agents.
@@
Diversity via Determinant (DvD) inspired by Parker-Holder et al. [27]. Calculate a policy behavior embedding and maximize the divergence in policy behavior as measured by the determinant of the behavioral embedding transformed by a kernel function.
@@ 4 Experiments
First, we describe our proposed benchmark MolExp, and show performance of various baseline RL algorithms with a comparison to an established benchmark. Then, we conducted three experiments to assess chemical exploration.
@@
    We investigated the effect of scaling TTT by the number of RL agents, or scaling TTT by the training time of a single RL agent by its molecule budget.
@@
    We proposed and tested different cooperative RL strategies to enhance exploration efficiency.
@@
    We translated our findings to a practically relevant task in optimizing the predicted bioactivity of generated compounds.
@@ 4.1 Molecule exploration benchmark
To address our proposed problem formulation to identify all hits in chemical space, we designed a new molecular exploration (MolExp) benchmark. Crucially, a ground truth is required to know that all solutions are identified  [^1] Oracles that are more practically relevant, such as bioactivity prediction or docking, don’t contain a ground truth set of all possible high scoring solutions.. For this reason, we use the toy-task of multiple molecule rediscovery via optimization of a similarity scoring function. To ensure relevance to real-world discovery, each set of molecules are at minimum pre-clinical drug candidates bioactive against the same target. Moreover, to make the objective more challenging, the set of molecules are selected to be structurally dissimilar albeit possessing similar bioactivity. This phenomena is the converse of the similarity principle  [^2] In drug discovery, the similarity principle states that similar molecules $Sim(m_{1},m_{2})\geq\epsilon$ possess similar bioactivity $|f(m_{1})-f(m_{2})|\leq\delta$. However, this does not prove the converse and many exceptions exist that are more difficult to model.. Note that activity cliffs are another example converse to the similarity principle that have been used to demonstrate more challenging scenarios for predictive models [39, 40].
@@
Each run is replicated on the benchmark 5 times with a default budget of 10,000 molecules per run, consistent with previous benchmarks [37].
Oracles
@@
The benchmark consists of four tasks. Each task is based on a set of drug candidates and their shared biological target, including anti-psychotic drugs (AP), Adenosine A2A (A2A) receptor drug candidates, beta-secretase 1 (BACE1) drug candidates, and epidermal growth factor drug candidates (EGFR). These tasks contain a set of 2, 3, 3, & 4 target molecules respectively, where the objective is to rediscover the full set of target molecules. To mimic a chemical space where distinct regions of reward signal exist, the oracle returns the similarity to the nearest target molecule as the reward $R(m)=max(sim(m_{i},m_{t_{1}}),...,sim(m_{i},m_{t_{N}}))$. As CLMs act in string space, we chose a string-edit similarity function based on Levenshtein distance (MolExpL). However, we additionally report the results when using fingerprint similarity (MolExp) in the Appendix. A more detailed description of the molecular targets chosen and the similarity function can be found in Appendix C.
Metric
@@
Performance is measured by calculating the product of the top N most similar molecules to each target molecule respectively $\prod_{t\in T}\frac{\sum_{i\in N}sim(m_{i},m_{t})}{N}$. This metric focuses on the ability generate similar molecules to all objectives. Unless stated otherwise, we use $N=1$ reflecting the maximum similarity achieved. Molecular diversity is also measured by sphere exclusion diversity (SEDiv) at a sample size of 1,000 [41] which provides a more accurate representation of chemical space coverage than internal diversity [42, 30].
Baseline performance
@@
We tested virtual screening, MolRL-MGPT [36], and multiple ACEGEN RL algorithms [23, 19] on their inherent ability for chemical exploration on the MolExpL benchmark (see Appendix D for implementation details). Note that MolRL-GPT was pre-trained on the same dataset for standardized comparison (see Appendix A for details). Table 1 shows that with a budget of 10,000 the ACEGENMolOpt algorithm achieved the highest score, therefore, we use this RL configuration for all further experiments. However, Figure 13 illustrates that none of the tested algorithms optimize all molecular targets within the default budget (for a single replicate), and that the best scores are due to improved efficiency in maximizing similarity to a single target molecule. Therefore, there is still considerable room for improvement. These results are also observed when using fingerprint similarity oracles (see subsection E.2).
Estimated max
@@
An estimated maximum performance can be calculated for ACEGENMolOpt RL configuration by training the agent to maximize only the similarity to each individual molecular target with the reward $R(m)=sim(m_{i},m_{t})$, (see Figure 14). This tests the agents ability to maximize similarity to each individual molecular target only, an expected prerequisite for maximizing similarity to multiple target molecules in the benchmark. Calculating the equivalent benchmark score results in an estimated maximum MolExpL score of 2.91.
Comparison to GuacaMol
For reference, baseline algorithms were tested on the GuacaMol benchmark also with a fixed budget of 10,000 molecules. This shows ACEGENMolOpt achieving the highest score (see subsection E.3). However, a key difference is that 5/6 GuacaMol similarity and rediscovery tasks are already solved by ACEGENMolOpt with a score $\geq 0.9$, whilst the MolExpL benchmark results in lower scores reflecting a higher difficulty of benchmark tasks.
Table 1: Baseline performance on the MolExpL benchmark. The ACEGENMolOpt RL configuration performs best in 3/4 tasks which we carry forward for scaling experiments.
@@
Task SCREEN MolRL-MGPT REINFORCE REINVENT REINVENTMolOpt AHC ACEGENPractical ACEGENMolOpt
AP 0.53 ± 0.04 0.48 ± 0.03 0.56 ± 0.05 0.57 ± 0.04 0.63 ± 0.02 0.62 ± 0.06 0.64 ± 0.02 0.64 ± 0.04
A2A 0.38 ± 0.03 0.37 ± 0.03 0.36 ± 0.02 0.39 ± 0.08 0.38 ± 0.03 0.40 ± 0.02 0.40 ± 0.07 0.44 ± 0.06
BACE1 0.29 ± 0.08 0.20 ± 0.02 0.23 ± 0.02 0.21 ± 0.01 0.27 ± 0.03 0.23 ± 0.01 0.23 ± 0.03 0.28 ± 0.03
EGFR 0.17 ± 0.04 0.17 ± 0.01 0.15 ± 0.02 0.16 ± 0.02 0.22 ± 0.03 0.17 ± 0.03 0.17 ± 0.02 0.25 ± 0.05
Sum 1.36 ± 0.10 1.21 ± 0.05 1.30 ± 0.06 1.33 ± 0.09 1.50 ± 0.06 1.41 ± 0.07 1.44 ± 0.08 1.62 ± 0.09
4.2 Test-time training scaling for chemical exploration